### v1

# Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b

11institutetext: Astrobiology Research Unit, Université de Liège, Allée du 6 Août 19C, B-4000 Liège, Belgium 22institutetext: Department of Earth, Atmospheric and Planetary Science, Massachusetts Institute of Technology, 77 Massachusetts Avenue, Cambridge, MA 02139, USA 33institutetext: Instituto de Astrofísica de Canarias (IAC), Calle Vía Láctea s/n, 38200, La Laguna, Tenerife, Spain 44institutetext: Center for Astrophysics |Harvard & Smithsonian, 60 Garden St, Cambridge, MA 02138, USA 55institutetext: Komaba Institute for Science, The University of Tokyo, 3-8-1 Komaba, Meguro, Tokyo 153-8902, Japan 66institutetext: Astrobiology Center, 2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan 77institutetext: Department of Physics, Sardar Vallabhbhai National Institute of Technology, Surat-395007, Gujarat, India 88institutetext: Dipartimento di Fisica ”Ettore Pancini”, Università di Napoli Federico II, Napoli, Italy 99institutetext: INFN, Sezione di Napoli, Complesso Universitario di Monte S. Angelo, Via Cintia Edificio 6, 80126 Napoli, Italy 1010institutetext: INAF - Osservatorio Astronomico di Capodimonte, via Moiariello 16, 80131 Napoli, Italy $1$institutetext: National Astronomical Observatory, 2-21-1 Osawa, Mitaka-shi, Tokyo 181-8588, Japan 1212institutetext: Department of Physics, Presidency University, Kolkata-700073, West Bengal, India 1313institutetext: Department of Astronomy & Astrophysics, University of California San Diego, La Jolla, CA 92093, USA 1414institutetext: Department of Physics and Kavli Institute for Astrophysics and Space Research, Massachusetts Institute of Technology, Cambridge, MA 02139, USA 1515institutetext: Departamento de Astrofísica, Universidad de La Laguna (ULL), E-38206 La Laguna, Tenerife, Spain 1616institutetext: Instituto de Astrofísica de Andalucía (IAA-CSIC), Glorieta de la Astronomía s/n, 18008 Granada, Spain 1717institutetext: Department of Physics and Astronomy, Vanderbilt University, Nashville, TN 37235, USA 1818institutetext: NSF’s National Optical-Infrared Astronomy Research Laboratory, 950 N. Cherry Ave., Tucson, AZ 85719, USA 1919institutetext: NASA Exoplanet Science Institute-Caltech/IPAC, Pasadena, CA 91125, USA 2020institutetext: Department of Physics, Center for Astrophysics & Space Sciences, University of California San Diego, La Jolla, CA 92093, USA 2121institutetext: Department of Physics & Astronomy, Stanford University, Stanford, CA 94305, USA $2$institutetext: Department of Multi-Disciplinary Sciences, Graduate School of Arts and Sciences, The University of Tokyo, 3-8-1 Komaba, Meguro, Tokyo 153-8902, Japan 2323institutetext: Dpt. Física Teórica y del Cosmos, Universidad de Granada, Campus de Fuentenueva s/n, 18071 Granada, Spain 2424institutetext: Department of Physics and Astronomy, University of Notre Dame, South Bend, IN 46556, USA 2525institutetext: NASA Ames Research Center, Moffett Field, CA 94035, USA 2626institutetext: Division of Science, National Astronomical Observatory of Japan, 2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan 2727institutetext: Astronomical Science Program, Graduate University for Advanced Studies, SOKENDAI, 2-21-1, Osawa, Mitaka, Tokyo, 181-8588, Japan 2828institutetext: Department of Physics and Kavli Institute for Astrophysics and Space Research, Massachusetts Institute of Technology, Cambridge, MA 02139, USA 2929institutetext: Center for Astrophysics and Space Sciences, UC San Diego, UCSD Mail Code 0424, 9500 Gilman Drive, La Jolla, CA 92093-0424, USA 3030institutetext: Department of Astronomy, University of Maryland, College Park, MD 20742, USA 3131institutetext: Kotizarovci Observatory, Sarsoni 90, 51216 Viskovo, Croatia 3232institutetext: Department of Astronomy, University of Tokyo, 7-3-1 Hongo, Bunkyo-ku, Tokyo 113-0033, Japan $3$institutetext: Department of Astrophysical and Planetary Sciences, University of Colorado Boulder, Boulder, CO 80309, USA 3434institutetext: Department of Astrophysical Sciences, Princeton University, Princeton, NJ 08544, USA 3535institutetext: SETI Institute, Mountain View, CA 94043 USA
Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b
K. Barkaoui ID E-mail: khalid.barkaoui@uliege.beThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    R. P. Schwarz ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    N. Narita ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    P. Mistry ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    C. Magliano ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    T. Hirano ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    M. Maity ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    A.J. Burgasser ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    B.V. Rackham ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    F. Murgas Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    F.J. Pozuelos ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    K. G. Stassun ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    M.E. Everett ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    D.R. Ciardi ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    C. Lamman ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    C. Aganze ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    E. Esparza-Borges ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    K.A. Collins ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    G. Covone ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    J. de Leon ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    M. Dévora-Pajares ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    J. de Wit Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    Izuru Fukuda ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    A. Fukui ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    R. Gerasimov ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    M. Gillon ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    Y. Hayashi ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    S.B. Howell ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    M. Ikoma ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    K. Ikuta ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    J.M. Jenkins ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    P.R. Karpoor ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    Y. Kawai ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    T. Kimura Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    T. Kotani Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    D.W. Latham ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    M. Mori ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    E. Pallé Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    H. Parviainen ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    Y.G. Patel ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    G. Ricker Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    H.M. Relles ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    A. Shporer ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    S. Seager ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    E. Softich ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    G. Srdoc Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    M. Tamura ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    C.A. Theissen ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    J.D. Twicken Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    R. Vanderspek Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    N. Watanabe ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    C.N. Watkins ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    J.N. Winn Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b    B. Wohler ID Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 bThree short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b
(Received/accepted)
One of the main goals of the NASA’s TESS (Transiting Exoplanet Survey Satellite) mission is the discovery of Earth-like planets around nearby M-dwarf stars. Here, we present the discovery and validation of three new short-period Earth-sized planets orbiting nearby M-dwarfs: TOI-5720 b, TOI-6008 b and TOI-6086 b. We combined TESS data, ground-based multi-color light curves, ground-based optical and near-infrared spectroscopy, and Subaru/IRD RVs data to validate the planetary candidates and constrain the physical parameters of the systems. In addition, we used archival images, high-resolution imaging, and statistical validation techniques to support the planetary validation. TOI-5720 b is a planet with a radius of $R_{p}=1.09\pm 0.07\leavevmode\nobreak\ R_{\oplus}$ orbiting a nearby (23 pc) M2.5 host, with an orbital period of $P=1.4344555\pm 0.0000036$ days. It has an equilibrium temperature of $T_{\rm eq}=708\pm 19$ K (assuming a null albedo) and an incident flux of $S_{p}=41.7\pm 4.5\leavevmode\nobreak\ S_{\oplus}$. TOI-6008 b has a period of $P=0.8574347\pm 0.0000424$ day, a radius of $R_{p}=1.03\pm 0.05\leavevmode\nobreak\ R_{\oplus}$, an equilibrium temperature of $T_{\rm eq}=707\pm 19$ K and an incident flux of $S_{p}=41.5\pm 4.5\leavevmode\nobreak\ S_{\oplus}$. The host star (TOI-6008) is a nearby (36 pc) M5 with an effective temperature of $T_{\rm eff}=3075\pm 75\leavevmode\nobreak\ K$. Based on the RV measurements collected with Subaru/IRD, we set a $3\sigma$ upper limit of $M_{p}<4M_{\oplus}$, thus ruling out a star or brown dwarf as the transiting companion. TOI-6086 b orbits its nearby (31 pc) M3 host star ($T_{\rm eff}=3200\pm 75$ K) every $1.3888725\pm 0.0000827$ days, and has a radius of $R_{p}=1.18\pm 0.07\leavevmode\nobreak\ R_{\oplus}$, an equilibrium temperature of $T_{\rm eq}=634\pm 16$ K and an incident flux of $S_{p}=26.8\pm 2.7\leavevmode\nobreak\ S_{\oplus}$. Additional high precision radial velocity measurements are needed to derive the planetary masses and bulk densities, and to search for additional planets in the systems. Moreover, short-period earth-sized planets orbiting around nearby M-dwarfs are suitable targets for atmospheric characterization with the James Webb Space Telescope ( JWST) through transmission and emission spectroscopy, and phase curve photometry.
Key Words.:
Exoplanetary systems; stars: TOI-5720, TOI-6008 and TOI-6086; techniques: photometric, techniques: radial velocity
1 Introduction
M dwarfs are stars with effective temperatures $<3800$ K (e.g. Reid et al. (1995); Nutzman & Charbonneau (2008); Kaltenegger & Traub (2009); Winters et al. (2014)). They are the most common stars in our galaxy (Henry et al. 1994; Kirkpatrick et al. 1999). M-dwarf stars are excellent targets to search for transiting Earth-like planets. Such systems offer a rare opportunity for exploring their physical, orbital and atmospheric parameters, due to their low masses, small sizes and low luminosities. The combination of the high equilibrium temperature of the planet, low mass, small radius and infrared brightness of the host star, making the short-period planets around M dwarfs exciting targets for detailed atmospheric characterization using the emission and transmission spectroscopy, and phase curve photometry with JWST (Zieba et al. 2023; Greene et al. 2023). Those planets with especially short periods (of order one day) may provide us some additional information into the formation, evolution and interior structure of exoplanetary systems (Winn et al. 2018). Moreover, they are also excellent systems to derive the slope of the radius valley (see, e.g., Fulton et al. (2017)), and determine star-planet interactions, such as the tidal dissipation’ effect within their host stars (see, e.g., Aller et al. (2020)).
In this context, we present the discovery and validation of three short-period Earth-sized TESS planets around the nearby M dwarfs TOI-5720, TOI-6008 and TOI-6086. By combining TESS data, and ground-based photometric and spectroscopic measurements, we find that TOI-5720 b, TOI-6008 b and TOI-6086 b are $R_{p}=1.09\pm 0.07\leavevmode\nobreak\ R_{\oplus}$, $R_{p}=1.03\pm 0.05\leavevmode\nobreak\ R_{\oplus}$ and $R_{p}=1.18\pm 0.07\leavevmode\nobreak\ R_{\oplus}$ with orbital periods of $P=1.4344555\pm 0.0000036$, $P=0.8574347\pm 0.0000424$ and $P=1.3888725\pm 0.0000827$ days, respectively.
The paper is organized as follows. Section 2 describes the TESS and ground-based observations. Section 3 describes the stellar characterisation of the host-stars using spectral energy distributions (SEDs) and spectroscopic observations. Section 4 presents the validation of the transit signals in the light curves. The planet searches and detection limits from the TESS photometry are presented in Section 5. Section 6 presents our global analysis of all available photometric and spectroscopic datasets of the planetary systems, allowing us to derive the physical parameters of the systems. Our discussion and conclusion are presented in Section 7.
2 Observations and data reduction
2.1 TESS  photometry
The host stars TIC 230055368 (TOI-5720), TIC 286201103 (TOI-6008) and TIC 18318288 (TOI-6086) were observed by the TESS  (Ricker et al. 2015) mission with 2-minutes cadence. TOI-5720 was observed in sectors 22 (2020 Feb 18 to 2020 Mar 18) and 49 (2022 Feb 26 to 2022 Mar 26). TOI-6008 was observed in sectors 15 (2019 Aug 15 to 2019 Sep 11), 16 (2019 Sep 11 to 2019 Oct 07), 41 (2021 Jul 23 to 2021 Aug 20), 55 (2022 Aug 05 to 2022 Sep 01), 56 (2022 Sep 01 to 2022 Sep 30) and 75 (2024 Jan 30 to 2024 Feb 26). TOI-6086 was observed in sectors 25 (2020 May 13 to 2020 Jun 08), 26 (2020 Jun 08 to 2020 Jul 04), 52 (2022 May 18 to 2022 Jun 13) and 53 (2022 Jun 13 to 2022 Jul 09). To perform TESS data modeling, we retrieved the Presearch Data Conditioning light curves (PDC-SAP, Stumpe et al. (2012); Smith et al. (2012); Stumpe et al. (2014) constructed by the TESS Science Processing Operations Center (SPOC; Jenkins et al. (2016)) at Ames Research Center from the Mikulski Archive for Space Telescopes. The PDC-SAP light curves have been calibrated for any instrument systematics and crowding effects. The TESS field-of-view for each target and photometric apertures used with the location of nearby Gaia DR3 sources (Gaia Collaboration et al. 2021a) are shown in Figure 1. Figures 2, 3 and 16 show the TESS photometric data for each target.
Refer to captionRefer to captionRefer to caption
Figure 1: TESS target pixel file images of TOI-5720 (left panel) TOI-6008 (middle panel) and TOI-6086 (right panel) observed in Sectors 22, 15 and 25, respectively. The mages are made by the tpfplotter package (Aller et al. 2020). Red dots and red shaded polygons show the position of Gaia DR3 sources and photometric apertures used to extract the photometric measurements, respectively.
Refer to captionRefer to captionRefer to caption
Figure 2: TESS PDC-SAP flux extracted from the 2-minutes cadence data of TOI-5720 (top panel), TOI-6008 (middle panel) and TOI-6086 (bottom panel). TOI-5720 was observed in Sectors 22 and 49, TOI-6008 was observed in Sectors 15, 16, 41, 55, 56 and 75, and TOI-6086 was observed in Sectors 25, 26, 52 and 53. Light grey points show the 2-minutes cadence data, and the black points shows the flux in 30-minutes bins. The transit locations of TOI-5720 b, TOI-6008 b and TOI-6086 b are shown with the blue triangles.
Refer to captionRefer to captionRefer to caption
Figure 3: TESS and ground-based photometric data of TOI-5720 b (left panel), TOI-6008 b (middle panel) and TOI-6086 b (right panel). Top plots show the TESS phased-folded, and other plots show the ground-based phased-folded by filter with best-fitting transit model is superimposed. The light curves are shifted along the y-axis for clarity.
2.2 Ground-based photometric follow-up
We performed ground-based observations follow-up of TOI-5720, TOI-6008 and TOI-6086 as part of the TESS Follow-up Observing Program (TFOP) in order to (i) rule out the NEBs (Nearby Eclipsing Binaries) as potential sources of the TESS detection, (ii) confirm the transit event on the target star, (iii) check the chromaticity of the transit depth measurement, (iv) refine the transit ephemerides. We used the TESS Transit Finder tool (Jensen 2013) in order to schedule the photometric follow-up observations. All ground-based photometric observations are summarized in Table 1.
2.2.1 LCOGT-2.0m MuSCAT3
We used the LCOGT (Las Cumbres Observatory Global Telescope; Brown et al. (2013)) 2.0m Faulkes Telescope North located at Haleakala Observatory in Hawaii to observe TOI-5720 b, TOI-6008 b and TOI-6086 b. The telescope is equipped with the MuSCAT3 multi-band imager (Narita et al. 2020). One full transit of TOI-5720 b was observed on UTC May 12 2023, three full transits of TOI-6008 b were observed on UTC May 07, 13 and 31 2023, and two full transits of TOI-6086 b were observed on UTC July 06 and 31 2023. All observations were carried out simultaneously with the Sloan-$g^{\prime}$, -$r^{\prime}$, -$i^{\prime}$ and Pan-STARRS-$z_{\mathrm{s}}$ filters. The stellar fluxes were extracted using uncontaminated photometric apertures (see Table 1).
The science images were calibrated using the standard LCOGT BANZAI pipeline (McCully et al. 2018), and aperture photometry was performed using AstroImageJ [^1] AstroImageJ: https://www.astro.louisville.edu/software/astroimagej/ (Collins et al. 2017) software.
2.2.2 MuSCAT2
We observed one full transit of TOI-6008 b on UTC July 27 2023 and one full transit of TOI-6086 b on UTC July 29 2023 with the 1.52m Telescopio Carlos Sanchez (TCS) at the Teide Observatory in Spain. The telescope is equipped with the MuSCAT2 multicolor imager (Narita et al. 2019). Both of transits were observed simultaneously in the Sloan-$g$, -$r$, -$i$, and Pan-STARRS-$z_{\mathrm{s}}$. The stellar fluxes were extracted using an uncontaminated photometric apertures (see Table 1). The data calibration and aperture photometry were performed using the MuSCAT2 photometry pipeline (Parviainen et al. 2020).
2.2.3 LCOGT-1.0m
Two full transits of TOI-6086 b were observed with LCOGT-1m0 network on UTC March 10 and 20 2023 in the Sloan-$i^{\prime}$ filter. The telescopes are equipped with $4096\times 4096$ SINISTRO Cameras with a pixel scale of $0.389\arcsec$ per pixel and a FOV of $26^{\prime}\times 26^{\prime}$. The science images were calibrated by the standard LCOGT BANZAI pipeline and aperture photometry was performed using AstroImageJ software. Both of observations are used for checking nearby eclipsing binaries, but they are not included in our global analysis because of the low signal-to-noise ratio.
Planet Date (UT) Filter Telescope Exptime FWHM Aperture Baseline $\beta_{w}$ $\beta_{r}$ $CF$ Comment
[second] [arcsec] [arcsec]
May 12 2023 $z_{\mathrm{s}}$ MuSCAT3 17 3.3 4.5 $p(t^{2})$ 1.26 1.08 1.36 Full
TOI-5720 b May 12 2023 $i^{\prime}$ MuSCAT3 19 3.3 4.5 $p(t^{2}+b^{1})$ 1.22 1.04 1.27 Full
May 12 2023 $r^{\prime}$ MuSCAT3 25 2.9 4.5 $p(t^{2}+b^{1})$ 1.26 1.09 1.37 Full
May 12 2023 $g^{\prime}$ MuSCAT3 209 2.6 4.5 $p(t^{2}+a^{1})$ 0.92 1.07 0.99 Full
May 07 2023 $z_{\mathrm{s}}$ MuSCAT3 20 2.1 3.2 $p(t^{2})$ 1.13 1.45 1.64 Full
May 07 2023 $i^{\prime}$ MuSCAT3 30 2.2 3.2 $p(t^{2})$ 1.13 1.20 1.36 Full
May 07 2023 $r^{\prime}$ MuSCAT3 88 2.3 3.2 $p(t^{2})$ 1.24 1.22 1.52 Full
May 13 2023 $z_{\mathrm{s}}$ MuSCAT3 20 2.9 4.5 $p(t^{2}+b^{1})$ 1.17 1.23 1.44 Full
May 13 2023 $i^{\prime}$ MuSCAT3 30 3.5 4.8 $p(t^{2})$ 1.16 1.30 1.51 Full
May 13 2023 $r^{\prime}$ MuSCAT3 88 3.5 5.0 $p(t^{2}+a^{2})$ 1.11 1.02 1.13 Full
TOI-6008 b May 13 2023 $g^{\prime}$ MuSCAT3 500 3.4 5.0 $p(t^{2})$ 0.85 1.23 1.04 Full
May 31 2023 $z_{\mathrm{s}}$ MuSCAT3 20 2.6 4.0 $p(x^{2})$ 1.15 2.46 2.82 Full
May 31 2023 $i^{\prime}$ MuSCAT3 30 2.9 4.2 $p(b^{2})$ 1.09 1.51 1.64 Full
May 31 2023 $r^{\prime}$ MuSCAT3 88 2.7 4.2 $p(t^{2})$ 1.05 1.19 1.26 Full
July 27 2023 $z_{\mathrm{s}}$ MuSCAT2 30 2.1 4.3 $p(t^{2})$ 1.02 1.31 1.34 Full
July 27 2023 $i^{\prime}$ MuSCAT2 30 2.4 4.4 $p(t^{2})$ 1.01 1.08 1.09 Full
July 27 2023 $r^{\prime}$ MuSCAT2 88 2.5 4.5 $p(t^{2})$ 0.98 1.18 1.18 Full
July 27 2023 $g^{\prime}$ MuSCAT2 30 2.7 4.5 $p(t^{2})$ 1.00 1.04 1.04 Full
July 06 2023 $z_{\mathrm{s}}$ MuSCAT3 35 2.4 4.8 $p(t^{2})$ 0.97 1.00 0.97 Full
July 06 2023 $i^{\prime}$ MuSCAT3 43 2.3 4.8 $p(t^{2})$ 0.95 1.00 0.95 Full
July 06 2023 $r^{\prime}$ MuSCAT3 63 2.5 4.8 $p(t^{2}+b^{1})$ 1.04 1.00 1.04 Full
July 31 2023 $z_{\mathrm{s}}$ MuSCAT3 35 4.1 6.9 $p(t^{2})$ 1.33 1.22 1.49 Full
July 31 2023 $i^{\prime}$ MuSCAT3 43 4.3 6.9 $p(t^{2})$ 1.16 1.04 1.21 Full
TOI-6086 b July 31 2023 $r^{\prime}$ MuSCAT3 63 4.3 6.9 $p(t^{2})$ 1.01 1.14 1.15 Full
July 29 2023 $z_{\mathrm{s}}$ MuSCAT2 20 2.1 10.9 $p(t^{2})$ 0.95 1.00 0.95 Full
July 29 2023 $i^{\prime}$ MuSCAT2 20 2.1 10.9 $p(t^{2})$ 1.02 1.09 1.011 Full
July 29 2023 $r^{\prime}$ MuSCAT2 40 2.5 10.9 $p(t^{2})$ 1.00 1.09 1.09 Full
July 29 2023 $g^{\prime}$ MuSCAT2 25 2.8 10.9 $p(t^{2})$ 1.00 1.36 1.36 Full
Table 1: Table shows the observational parameters for TOI-5720 b, TOI-6008 b and TOI-6086 b: date of observation, filter used, telescope, exposure time(s), and FWHM of the point-spread function and photometric aperture. Analysis parameters: the selected baseline model, red noise $\beta_{r}$, white noise $\beta_{w}$ and $CF$ (Correction Factor). The baseline model is shown as a polynomial $p()$ of systematics parameters, with $t=$ time, $x=$ and $y=$ positions of the star on the detector, $a=$ airmass, $b=$ background, and $f=$ full-with at half maximum of the PSF.
2.2.4 Shane/Kast Optical Spectroscopy
[…]
Refer to captionRefer to captionRefer to caption
[…]
2.2.5 IRTF/SpeX Spectroscopy
We used the SpeX spectrograph (Rayner et al. 2003) on the 3.2m-NASA IRTF (Infrared Telescope Facility) to collect medium-resolution near-infrared spectra of TOI-6086 on UT 2023 May 04 and of TOI-5720 and TOI-6008 on UT 2023 June 05. Thin cirrus clouds were present on both nights. Seeing was ${\prime\prime}$0 on the May night and ${\prime\prime}$6 on the June night. On both nights, we used the short-wavelength cross-dispersed (SXD) mode with the $0\aas@@fstack{\prime\prime}3\times 15^{\prime\prime}$ slit aligned to the parallactic angle, providing spectra covering 0.80–2.42 $\mu$m with a resolving power of $R{\sim}2000$ and an average of 2.5 pixels per resolution element. Nodding in an ABBA pattern, we collected six integrations of TOI-5720 and TOI-6008 and eight integrations of TOI-6086 with integration times of 70.0 s, 70.0 s, and 149.7 s, respectively. After each science observation, we collected a set of standard SXD flat-field and arc-lamp exposures and six integrations of A0 V standard stars. We reduced the data with Spextool v4.1 (Cushing et al. 2004), following the standard approach outlined in the Spextool User’s Manual [^3] Available at http://irtfweb.ifa.hawaii.edu/~spex/observer/.. The final spectra have median signal-to-noise ratios per pixel of 124, 109, and 133 for TOI-5720, TOI-6008, and TOI-6086, respectively (see Figure 5).
Figure 5: SpeX SXD spectra of TOI-5720 (top), TOI-6008 (middle), and TOI-6086 (bottom). The target spectra (blue) are shown alongside the spectrum of the best-fit standard star (grey). Regions of high telluric absorption are shaded, and strong spectral features of M dwarfs are highlighted.
2.3 Spectroscopy
2.3.1 Spectroscopy from Subaru/IRD
We performed high-resolution spectroscopy for TOI-6008 with the InfraRed Doppler (IRD) instrument (Tamura et al. 2012; Kotani et al. 2018) mounted at the Subaru 8.2m telescope. IRD can obtain a near-infrared spectrum between from 950 nm to 1740 nm at one exposure with a spectral resolution of $R\approx 70,000$. Between UT 2023 July 31 and August 6, a total of 15 frames were collected for TOI-6008, for which integration times were set to 1200–1500 seconds. We also injected the laser-frequency comb (LFC) into a secondary fiber so that we can trace the temporal variation of the instrumental profile.
IRD data were reduced following the reduction procedure in Hirano et al. (2020), where custom reduction codes as well as IRAF (Tody 1993) were used. For each frame, one-dimensional stellar and LFC spectra were separately extracted by this reduction. The reduced stellar spectra typically have signal-to-noise (S/N) ratios of 60–70 per pixel at 1000 nm. Those spectra were then subjected to the radial-velocity (RV) measurements with the standard RV pipeline for IRD (Hirano et al. 2020), and we extracted relative RVs for individual frames using the template-matching technique. The resulting RV measurements and curve are presented in Table 2 and Figure 6, respectively; the RV internal errors returned by the pipeline were typically $\approx 3$ m s-1.
BJD-TDB RV [km/s] $\sigma_{RV}$ [km/s]
2460156.9444130 6.337714 0.00332
2460156.9586420 6.344117 0.00310
2460156.9728711 6.332584 0.00301
2460159.9418833 6.345601 0.00327
2460159.9561104 6.347892 0.00330
2460159.9703249 6.343831 0.00323
2460160.9039142 6.342193 0.00282
2460160.9181436 6.338026 0.00282
2460160.9323684 6.339218 0.00286
2460161.8541920 6.335824 0.00348
2460161.8684328 6.342417 0.00321
2460161.8826727 6.339555 0.00325
2460162.8982452 6.334722 0.00306
2460162.9124742 6.333475 0.00303
2460162.9294896 6.337795 0.00274
Table 2: Radial-velocities measurements for TOI-6008 obtained from Subar-8.2m/IRD spectrograph.
Figure 6: RVs obtained with the Subaru-8.2m/IRD spectrograph for TOI-6008. a, RVs observations versus time. b, RVs observations phase-folded. The best-fitting model is superimposed in red.
2.3.2 Spectroscopy from TRES
Two reconnaissance spectra of TOI-5720 were obtained on UT December 19 and 24, 2022 using the Tillinghast Reflector Echelle Spectrograph (TRES; Fűrész (2008)) mounted on the 1.5m Tillinghast Reflector telescope at the Fred Lawrence Whipple Observatory (FLWO) in Arizona. TRES is a fiber-fed spectrograph that observes in the 390–910 nm range with a resolving power of R=44,000. Spectra were extracted using methods described in Buchhave et al. (2010).
For the two TRES observations of TOI-5720, we measure radial velocities using a pipeline optimized for mid-to-late M dwarfs (see e.g. Pass et al. (2023) for further details). In brief, this pipeline performs a cross correlation between the observed spectra and an empirical mid-to-late M dwarf template using six TRES echelle orders in the range of 6400–7850 Å. The pipeline also produces carefully calibrated radial velocity uncertainties, taking into account the signal-to-noise ratio of the observed spectra, template mismatch, rotational broadening as well as the long-term stability of the spectrograph and errors in the barycentric correction. The two resulting RVs measurements are $35.692\pm 0.031$ km/s on UT December 19 and $35.684\pm 0.024$ km/s for UT December 24, and an upper limit on the projected rotational velocity is found to be $v\sin(i)<3.4$ km/s.
2.4 High-Resolution imaging
We assess the possibility of a blend of nearby stellar companions as part of our standard process for validating transiting exoplanets, that can dilute the transit light curves and underestimated planetary radii measurements (Ciardi et al. 2015). We observed TOI-5720, TOI-6008 and TOI-6086 with WIYN 3.5 m/NESSI, Palomar-5m0/PHARO and Robo-AO high-resolution imaging. Despite the combination of Gaia catalog and high-resolution imaging, there is no indication of any additional close stellar companions.
2.4.1 Palomar-5m0/PHARO
We observed TOI-5720, TOI-6008 and TOI-6086 with the PHARO instrument (Hayward et al. 2001) behind the natural guide star AO system P3K (Dekany et al. 2013) on UT 2023 June 07 and 08 in the narrow-band $Br-\gamma$ filter ($\lambda_{0}=2.1686;\Delta\lambda=0.0326$ $\mu$m). PHARO equipped with a 1024$\times$1024 HgCdTe HAWAII detector, with a pixel scale of 0.025 arcsec, resulting a total FOV of 25 arcsec. Each dither position was observed three times for a total of 15 frames; with an exposure time of 9.9s for TOI-5720, 2.8s for TOI-6008 and 9.9s for TOI-6086. The total on-source time was 148.5s for TOI-5720, 42.5s for TOI-6008 and 148.5s for TOI-6086. The science data were calibrated by sky-subtracted and flat-fielded. Further details of observations from the Palomar-5m0/PHARO are available in Hayward et al. (2001). No companion sources were detected for TOI-5720, TOI-6008 and TOI-6086 in the PHARO data (see Figure 7).
Figure 7: High resolution imaging of TOI-5720 (top panel), TOI-6008 (middle pane), and TOI-6086 (bottom panel) obtained with the Palomar/PHARO Speckle Imager on UT 2023 June 7 and 8. No close stellar companion sources were detected near the targets.
3 Stellar characterisation
Star information
Parameter TOI-5720 TOI-6008 TOI-6086 Source
Identifying information:
TIC 230055368 286201103 18318288
GAIA DR3 ID 3994047048130880384 2168280502430898944 4597447518944594176
2MASS ID J11211567+2516255 J20414744+4938482 J17412905+3018135
Parallax and distance:
RA [J2000] 11:21:15.26 20:41:47.70 17:41:28.90 (1)
Dec [J2000] +25:16:23.86 +49:38:48.84 +30:18:13.59 (1)
Plx [$mas$] $27.878\pm 0.022$ $43.44\pm 0.016$ $31.7665\pm 0.0142$ (1)
$\mu_{RA}$ [mas yr-1] $-350.367\pm 0.115$ $158.442\pm 0.102$ $-110.595\pm 0.061$ (1)
$\mu_{Dec}$ [mas yr-1] $-91.492\pm 0.116$ $44.404\pm 0.093$ $2.165\pm 0.071$ (1)
Distance [pc] $35.87\pm 0.03$ $23.02\pm 0.01$ $31.47\pm 0.014$ (1)
Photometric properties:
TESSmag $11.619\pm 0.007$ $12.294\pm 0.008$ $12.411\pm 0.007$ (2)
$V_{\rm mag}$ [UCAC4] $13.847\pm 0.043$ $15.97\pm 0.20$ $14.98\pm 0.06$ (3)
$B_{\rm mag}$ [UCAC4] $15.469\pm 0.083$ $16.824$ $16.78\pm 0.06$ (3)
$J_{\rm mag}$ [2MASS] $10.123\pm 0.022$ $10.416\pm 0.024$ $10.82\pm 0.02$ (4)
$H_{\rm mag}$ [2MASS] $9.485\pm 0.019$ $9.830\pm 0.030$ $10.25\pm 0.02$ (4)
$K_{\rm mag}$ [2MASS] $9.288\pm 0.015$ $9.541\pm 0.018$ $9.99\pm 0.02$ (4)
$G_{\rm mag}$ [Gaia DR3] $15.594\pm 0.001$ $13.712\pm 0.001$ $13.797\pm 0.001$ (1)
$W1_{\rm mag}$ [WISE] $9.134\pm 0.023$ $9.334\pm 0.022$ $9.80\pm 0.02$ (5)
$W2_{\rm mag}$ [WISE] $8.988\pm 0.020$ $9.183\pm 0.02$ $9.62\pm 0.02$ (5)
$W3_{\rm mag}$ [WISE] $8.892\pm 0.029$ $9.234\pm 0.053$ $9.53\pm 0.04$ (5)
$W4_{\rm mag}$ [WISE] $8.835\pm 0.438$ $9.128$ $8.87\pm 0.46$ (5)
Spectroscopic and derived parameters
Optical spectral type M3.5 $\pm$ 1.0 M5.0 M3.0 this work
Near-infrared spectral type M3.5 $\pm$ 1.0 M5.0 M3.5 this work
SED $T_{\rm eff}$ [K] $3325\pm 75$ $3075\pm 75$ $3200\pm 75$ this work
SED $\log g_{\star}$ [dex] $5.01\pm 0.04$ $5.01\pm 0.02$ $5.01\pm 0.02$ this work
Optical [Fe/H] [dex] $+0.03\pm 0.20$ $\gtrsim+0.3$ $+0.11\pm 0.20$ this work
Near-infrared [Fe/H] [dex] $-0.19\pm 0.12$ $+0.21\pm 0.12$ $-0.20\pm 0.12$ this work
SED [Fe/H] [dex] $0.0\pm 0.3$ $+0.3\pm 0.3$ $-0.2\pm 0.3$ this work
$v\sin(i)$ [km/s] $<3.4$ this work
$M_{\star}$ [$M_{\odot}$] $0.383\pm 0.019$ $0.230\pm 0.011$ $0.254\pm 0.013$ this work
$R_{\star}$ [$R_{\odot}$] $0.383\pm 0.018$ $0.242\pm 0.013$ $0.259\pm 0.013$ this work
$\rho_{\star}$ [$\rho_{\odot}$] $6.82\pm 1.02$ $16.22\pm 2.73$ $14.62\pm 2.3$ this work
$F_{\rm bol}\times 10^{-10}$ [erg s-1 cm-2] $4.027\pm 0.094$ $2.839\pm 0.101$ $2.051\pm 0.072$ this work
$L_{\rm bol}$ [L] $0.01615\pm 0.00019$ $0.004691\pm 0.000082$ $0.00634\pm 0.00011$ this work
$Av$ [mag] $0.00\pm 0.02$ $0.00\pm 0.02$ $0.00\pm 0.02$ this work
[…]
Table 3: Spectroscopy, astrometry and photometry stellar properties of the host stars TOI-5720, TOI-6008 and TOI-6086. (1) Gaia EDR3; Gaia Collaboration et al. (2021a); (2) TESS Input Catalog; Stassun et al. (2018b); (3) UCAC4; Zacharias et al. (2012); (4) 2MASS; Skrutskie et al. (2006); (5) WISE; Cutri et al. (2021).
Parameter TOI-5720 b TOI-6008 b TOI-6086 b
Planet/star area ratio
$(R_{p}/R_{\star})^{2}_{\rm TESS}$ [ppm] $725^{+188}_{-186}$ $1517^{-120}_{+122}$ $1531^{-370}_{+376}$
$(R_{p}/R_{\star})^{2}_{{\rm Sloan}-z^{\prime}}$ [ppm] $989^{+156}_{-155}$ $1431^{-150}_{+151}$ $1783^{-261}_{+269}$
$(R_{p}/R_{\star})^{2}_{{\rm Sloan}-i^{\prime}}$ [ppm] $723^{+193}_{-190}$ $1591^{-186}_{+188}$ $1969^{-275}_{+286}$
$(R_{p}/R_{\star})^{2}_{{\rm Sloan}-r^{\prime}}$ [ppm] $635^{+218}_{-223}$ $1420^{-292}_{+301}$ $1639^{-401}_{+405}$
$(R_{p}/R_{\star})^{2}_{{\rm Sloan}-g^{\prime}}$ [ppm] $724^{+188}_{-186}$ $970^{-610}_{+746}$ $2951^{-1260}_{+1902}$
Orbital period $P$ [days] $1.4344555\pm 0.0000036$ $0.8574347\pm 0.0000424$ $1.3888725\pm 0.0000827$
Transit-timing $T_{0}$ $10076.902501\pm 0.00067$ $10078.016214\pm 0.000311$ $10131.969501\pm 0.000567$
$[BJD_{\rm TDB}-2450000]$
Orbital semi-major axis $a$ [AU] $0.018288\pm 0.000279$ $0.010777\pm 0.000162$ $0.01541\pm 0.00024$
Impact parameter $b$ [$R_{\star}$] $0.14\pm 0.09$ $0.45\pm 0.08$ $0.608\pm 0.051$
Transit duration $W$ [min] $61\pm 2$ $39\pm 1$ $42\pm 2$
Scaled semi-major axis $a/R_{\star}$ $10.97\pm 0.29$ $9.44\pm 0.21$ $12.75\pm 0.26$
Orbital inclination $i$ [deg] $89.3\pm 0.6$ $87.2\pm 0.5$ $87.3\pm 0.3$
Radius ratio $R_{\oplus}/R_{\star}$ $0.02796\pm 0.00158$ $0.03867\pm 0.00109$ $0.04173\pm 0.00220$
Radius $R_{p}$ [$R_{\oplus}$] $1.09\pm 0.07$ $1.03\pm 0.05$ $1.18\pm 0.07$
Irradiation $S_{p}$ [$S_{\oplus}$] $41.7\pm 4.5$ $41.5\pm 4.5$ $26.8\pm 2.7$
a Equilibrium temperature $T_{\rm eq}$ [K] $708\pm 19$ $707\pm 19$ $634\pm 16$
b Mass $M_{p}$ [$M_{\oplus}$ ] $<4.3$ $1.1^{+3.3}_{-0.9}$
b RV semi-amplitude $K$ [$m/s$] $<64$ $2.1^{+5.9}_{-1.8}$
c Predicted mass $M_{p}$ [$M_{\oplus}$ ] $1.32^{+0.92}_{-0.47}$ $1.10^{+0.73}_{-0.36}$ $1.68^{+1.15}_{-0.56}$
c Predicted RV semi-amplitude $K$ [$m/s$] $1.4^{+1.0}_{-0.5}$ $2.0^{+1.3}_{-0.6}$ $2.4^{+1.6}_{-0.8}$
d Predicted TSM $9.7^{+7.1}_{-4.1}$ $18.8^{+12.6}_{-6.8}$ $12.0^{+8.5}_{-4.7}$
d Predicted ESM $3.4\pm 0.5$ 6.4$\pm$0.7 4.2$\pm$0.7
Table 4: Planetary physical parameters determined from our global MCMC analysis (medians and $1\sigma$ error bar) for TOI-5720 b, TOI-6008 b and TOI-6086 b. a The equilibrium temperature derived assuming a null albedo. b Planetary mass and RV semi-amplitude values derived from RVs measurements collected by Subaru/IRD and TrES spectrographs for TOI-6008 b and TOI-5720 b, respectively. c Predicted mass and RV semi-amplitude values calculated using the mass-radius relationship from Chen & Kipping (2017). d TSM (Transmission spectroscopy metric) and ESM (emission spectroscopy metric) values calculated using equations 1 and 4 from Kempton et al. (2018).
Parameter TOI-5720 TOI-6008 TOI-6086
Mean density, $\rho_{\star}$ [$\rho_{\odot}$] $8.64^{+0.77}_{-0.67}$ $15.41^{+1.02}_{-0.95}$ $14.41^{+0.96}_{-0.89}$
Stellar mass, $M_{\star}$ [$M_{\odot}$ ] $0.397^{+0.017}_{-0.018}$ $0.227^{+0.010}_{-0.011}$ $0.253^{+0.012}_{-0.012}$
Stellar radius, $R_{\star}$ [$R_{\odot}$] $0.358^{+0.011}_{-0.011}$ $0.245^{+0.007}_{-0.007}$ $0.260^{+0.007}_{-0.008}$
Luminosity, $L_{\star}$ [$L_{\odot}$] $0.01402_{-0.00145}^{+0.00162}$ $0.00480^{+0.00058}_{-0.00053}$ $0.00639^{+0.00077}_{-0.00068}$
Effective temperature, $T_{\rm eff}$ [K] $3324^{+73}_{-76}$ $3073^{+75}_{-76}$ $3205^{+76}_{-77}$
Quadratic Limb-Darkening coefficients
$u_{\rm 1,TESS}$ $0.26\pm 0.01$ $0.41\pm 0.03$ $0.21\pm 0.02$
$u_{\rm 2,TESS}$ $0.39\pm 0.01$ $0.14\pm 0.01$ $0.42\pm 0.01$
$u_{\rm 1,{\rm Pan-STARRS-z_{s}}}$ $0.16\pm 0.01$ $0.34\pm 0.02$ $0.15\pm 0.01$
$u_{\rm 2,{\rm Pan-STARRS-z_{s}}}$ $0.42\pm 0.01$ $0.13\pm 0.01$ $0.43\pm 0.01$
$u_{\rm 1,{\rm Sloan}-i^{\prime}}$ $0.30\pm 0.01$ $0.42\pm 0.02$ $0.29\pm 0.01$
$u_{\rm 2,{\rm Sloan}-i^{\prime}}$ $0.35\pm 0.01$ $0.22\pm 0.01$ $0.36\pm 0.01$
$u_{\rm 1,{\rm Sloan}-r^{\prime}}$ $0.42\pm 0.01$ $0.57\pm 0.02$ $0.40\pm 0.01$
$u_{\rm 2,{\rm Sloan}-r^{\prime}}$ $0.34\pm 0.01$ $0.26\pm 0.01$ $0.36\pm 0.01$
$u_{\rm 1,{\rm Sloan}-g^{\prime}}$ $0.41\pm 0.01$ $0.61\pm 0.02$ $0.39\pm 0.02$
$u_{\rm 2,{\rm Sloan}-g^{\prime}}$ $0.37\pm 0.01$ $0.26\pm 0.01$ $0.40\pm 0.04$
Table 5: Derived stellar parameters ($\rho_{\star}$, $M_{\star}$, $R_{\star}$, $L_{\star}$, $T_{\rm eff}$) and Limb-darkening coefficients for TOI-5720, TOI-6008 and TOI-6086 based on our global analysis.
3.1 Shane/Kast spectroscopic analysis
[…]
3.2 IRTF/SpeX spectroscopic analysis
The SpeX SXD spectra of TOI-5720, TOI-6008, and TOI-6086 are shown in Figure 5. Using the SpeX Prism Library Analysis Toolkit (SPLAT, Burgasser & Splat Development Team 2017), we assigned spectral types by comparing the spectra to those of single-star spectral standards in the IRTF Spectral Library (Cushing et al. 2005; Rayner et al. 2009). Based upon the best matches, we adopt spectral types of M2.5 $\pm$ 0.5, M5.0 $\pm$ 0.5, and M3.5 $\pm$ 0.5 for TOI-5720, TOI-6008, and TOI-6086, respectively. We also used SPLAT to measure the equivalent widths of the $K$-band Na i and Ca i doublets and the H2O–K2 index (Rojas-Ayala et al. 2012). Using the Mann et al. (2013b) relation between these observables and stellar metallicity, and propagating uncertainties using a Monte Carlo approach (see Delrez et al. 2022), we estimate iron abundances of $\mathrm{[Fe/H]}=-0.19\pm 0.12$ for TOI-5720, $\mathrm{[Fe/H]}=+0.21\pm 0.12$ for TOI-6008, and $\mathrm{[Fe/H]}=-0.20\pm 0.12$ for TOI-6086 (see Table 3).
3.3 SED analysis and evolutionary models
We conducted an independent analysis of the broadband SED (spectral energy distribution) of the stars together with the Gaia DR3 parallax (with no systematic offset applied; see, e.g., Stassun & Torres 2021), to derive an empirical measurement of the stellar radii, following the same method described in Stassun & Torres (2016); Stassun et al. (2017, 2018a). We pulled the 2MASS $JHK_{S}$ magnitudes, the WISE W1–W4 magnitudes, the Pan-STARRS $grizy$ magnitudes, as well as the Gaia $G_{\rm BP}G_{\rm RP}$ magnitudes. Together, the available photometry spans the full stellar SED over the wavelength range 0.4–20 $\mu$m (see Figure 8).
Figure 8: Spectral Energy Distribution (SED) fit of TOI-5720 (left), TOI-6008 (middle) and TOI-6086 (right). The gray curves are the best-fitting PHOENIX atmosphere model, colored symbols with error-bars are the observed fluxes (horizontal bars represent the effective bandpasses), and black symbols are the model fluxes.
We performed a fit using PHOENIX stellar atmosphere models (Husser et al. 2013a), with the free parameters being the metallicity ([Fe/H]) and effective temperature ($T_{\rm eff}$) we set null extinction due to the close proximity of the stars to Earth. The resulting fits (Figure 8) have a best-fit $T_{\rm eff}=3325\pm 75$ K, $T_{\rm eff}=3075\pm 75$ K, and $T_{\rm eff}=3200\pm 75$ K for TOI-5720, TOI-6008, and TOI-6086 respectively, and a best-fit [Fe/H] = $0.0\pm 0.3$, $+0.3\pm 0.3$, and $-0.2\pm 0.3$ respectively, with a reduced $\chi^{2}$ of 1.8, 1.9, and 2.4 respectively (the latter star exhibits some UV excess that may indicate chromospheric activity). Integrating the model SED gives the bolometric flux at Earth, $F_{\rm bol}=4.027\pm 0.094\times 10^{-10}$ erg s-1 cm-2, $F_{\rm bol}=2.839\pm 0.101\times 10^{-10}$ erg s-1 cm-2, and $F_{\rm bol}=2.051\pm 0.072\times 10^{-10}$ erg s-1 cm-2 respectively. Taking the $F_{\rm bol}$ together with the Gaia parallax, gives the bolometric luminosity directly as $L_{\rm bol}=0.01615\pm 0.00019$ L, $L_{\rm bol}=0.004691\pm 0.000082$ L, and $L_{\rm bol}=0.00634\pm 0.00011$ L respectively. The stellar radii then follow from the Stefan-Boltzmann relation as $R_{\star}=0.383\pm 0.018$ R, $R_{\star}=0.242\pm 0.013$ R, and $R_{\star}=0.259\pm 0.013$ R respectively. Finally, we can estimate the stellar mass from the empirical mass-luminosity relation of Mann et al. (2019), giving $M_{\star}=0.383\pm 0.019$ M, $M_{\star}=0.230\pm 0.011$ M, and $M_{\star}=0.254\pm 0.013$ M respectively.
4 Planet validation
4.1 Data validation report for TESS
As we discribed in Section 2.1, the SPOC pipeline extracted the photometric measurements for TOI-5720, TOI-6008 and TOI-6086 using 2-minutes cadence, yielding orbital periods of $P=1.43446$ days, $P=0.85737$ day, and $P=1.38886$ days, and signal-to-noise ratios of $S/N=7.7$, $S/N=14.9$ and $S/N=8.1$, respectively. The TOI vetting team reviewed the SPOC Data Validation reports (Jenkins 2002; Jenkins et al. 2010, 2020) on May 26 2022, Sept 16 2022 and Sept 15 2022 (Guerrero et al. 2021) for TOI-5720, TOI-6008 and TOI-6086, respectively. The transit depths were found to be $1.03\pm 0.16$ ppt (for TOI-5720.01), $2.05\pm 0.27$ ppt (for TOI-6008.01) and $1.37\pm 0.24$ ppt (for TOI-6086.01), corresponding to a planet radius of $R_{p}=1.4\pm 0.3\pm$ R, $R_{p}=1.2\pm 0.2\pm$ R and $1.0\pm 1.7\pm$ R for TOI-5720.01, TOI-6008.01 and TOI-6086.01, respectively. All additional validation tests, including bootstrap, centroid offset, ghost, difference imaging centroid tests, as well as a search for discrepancies between odd and even transit depths were successfully passed. These conclusive tests prompted us to conduct further high-precision photometric follow-up and analysis to validate the planet candidates.
4.2 Vetting
While the transit method is one of the most fruitful technique for discovering exoplanets, several scenarios other than a transiting exoplanet can produce periodic dips in the brightness of a star. Numerous false positive scenarios exist in astronomical observations: i) stellar binary systems, where two stars sharing a common center might create a signal resembling a planetary transit if their sizes differ or if their eclipse is partial; ii) blended binary stars, wherein the target blends the light of an eclipsing binary obscuring the secondary eclipse and confusing the primary with a planetary transit; iii) other misleading signals from astrophysical sources (stellar spots, pulsations, rotation) or instrument-related factors (jitter noise, momentum dumps).
Thus, the first essential step towards the confirmation of a transiting exoplanetary signal is the vetting. In addition to the SPOC DV reports for TOI 5720.01, TOI 6008.01 and TOI 6086.01, we conduct a uniform vetting analysis through DAVE (Discovery and Vetting of Exoplanets, Kostov et al. 2019) which allows to examine a transit event on a double-level analysis: the pixel level by a photocenter analysis and the light curve level by a flux time-series analysis. The centroid module generates a difference image by subtracting the overall in-transit image from the out-of-transit image. It determines the photocenter by fitting the TESS pixel response function to the image, performed for each detected transit. The average centroid position across events is calculated, along with its statistical significance, aiding the vetting process. An offset in the centroid position from the expected target star’s position flags the event as a False Positive (FP). However, challenging interpretation due to artifacts or low SNR in DAVE-produced difference images makes the photocenter analysis unreliable. In such cases we classify the signal as a planet candidate. The Modelshift module produces a phase-folded light curve with the best-fit trapezoid transit model. Its primary goal is to determine if the signal source is an eclipsing binary system. It showcases key transit features, including the average primary signal and other significant features. Additionally, it compares odd and even transits, measuring their statistical difference. It also evaluates the transit shape, offering a comprehensive assessment of the observed transit properties.
In the vetting analysis of planetary transit candidates, reliability can be compromised when the SNR of the event is particularly low. This occurs because low SNR signals may be indistinguishable from noise or other sources of variability in the data, making it challenging to accurately identify genuine transit events. As a result, vetting efforts may be less effective in confirming or ruling out potential candidates in such cases. TOI 5720.01, TOI 6008.01 and TOI 6086.01 are three low SNR planet candidates with the first and the third being near the detection threshold (SNR=$7.1$, Borucki et al. 2011). However their transit depths on the phase-folded light curve are slightly above the noise level. Moreover, we did not flag any particular anomalies in the shape of the transit as well as no statistical significant evidence arose between the odd and the even transits. As the centroid module is concerned, TOI 5720.01 passed the photocenter test without any concern. The same test was inconclusive for both TOI 6008.01 and TOI 6086.01 since the TESS aperture masks used to extract their light curves are crowded with contaminating stars, resulting in an unreliable overall centroid measurement. However, for a few single clean centroid measurements the light photocenters are consistent with the position of both TOI 6008 and TOI 6086, as we will discuss in Sect.4.3.
4.3 Ancillary information
Since the three planet candidates have a low SNR, we conducted further analysis other than those provided by DAVE to make our vetting process more solid. TESS has a large pixel scale, about $21^{\prime\prime}$/pixel, with a focus-limited PSF. In order to check whether any resolved or unresolved sources fell within the aperture mask, thus contaminating the photometric signal, we consulted stellar catalogs (Wenger et al. 2000; Gaia Collaboration et al. 2021b). While the centroids module should shed light on the contamination flux due to resolved nearby stars, the same cannot be said for the unresolved sources within the same pixel of the star target. At this level we can only flag the presence of a potentially contaminating source which dilutes the transit depth. The worst scenario is if the contaminating source is the origin of the transit event. We here stress this is an intrinsic limit of this analysis which can be overcome by follow-up photometric campaigns (see Sect. 4.6).
TOI 5720 has only a very faint (TESS magnitude $=18.162$) nearby target, TIC 903650250, which is $\approx 55$ arcsec apart and seven magnitudes dimmer.
On the other hand, TOI 6008 is located in a region of the sky overdense of stars ($82$ known targets in $1$ arcmin). In particular, $5$ stars fall within the same pixel of TOI 6008 even if the brightest one (TIC 286201079) is almost $5$ magnitudes fainter than TOI 6008. Moreover, the brightest star is TIC 286201082 (TESS magnitude $=11.57$) which is $41$ arcsec ($\approx 2$ pixels) away from our target and $\approx 1$ magnitude brighter than it. Thus, the centroid measurements are triggered by the photometric variability of TIC 286201082. Another $42$ arcsec distant star, TIC 286201047 (TESS magnitude $=13$), contributes to make the photocenter measurements challenging. Despite most centroid measurements are unreliable, a few are not altered by nearby sources. For those, the centroid is aligned with the TOI 6008’s position.
The photometry obtained from TOI 6086 observations are contaminated by the star light of TIC 18318284 (TESS magnitude $=13.25$) which is only one magnitude fainter than TOI 6086 and $46$ arcsec distant. Moreover, the photocenter measurements are completely counterfeit due to the presence of a $120^{\prime\prime}$ arcsec star (Gaia DR3 4597447931261454080) which is 2 magnitudes brighter than TOI 6086. Also in this case, a few clean centroid measurements do not show any flag for an offset.
We also decided to conduct a Pixel Level light curve (PLL) analysis using the publicly available LATTE [^4] LATTE: https://github.com/noraeisner/LATTE pipeline developed by Eisner (2022). This test allowed us to inspect the light curve for each pixel in the field of view of the corresponding target pixel file to check whether the transit occurs in nearby pixels of the mask. None of the pixels within the aperture mask other than that corresponding to TOI 6008 and TOI 6086 shows a reliable dip in the starlight. We repeated the same procedure for all the transits observed by TESS with the same conclusions.
Despite the orbital path of TESS has been designed to maximize the sky coverage while reducing the number of obstructions, the TESS Full Frame Images can occasionally be contaminated primarily by the zodiacal light and scattered light from solar system objects (Gangestad et al. 2013; Sullivan et al. 2015). Thus, during each TESS sector’s observation, the flux of the background can vary widely. If any unusual events in the background occur at the transit time of the planet candidate, they could introduce stilted signals into the light curve, contaminating or, even worse, originating the detected transit. In order to ensure that no anomaly contaminated the signal, we visually inspected a 3-day-long section of the background flux centered at the time of transit. We found no evidences of background artifacts for TOI 5720.01, TOI 6008.01 and TOI 6086.01.
We found no known TESS momentum dumps in correspondence of the transits of TOI 5720.01, TOI 6008.01 and TOI 6086.01, thus ruling out any possible instrumental systematic scenario.
4.4 Archival imaging
We used archival images of TOI-5720, TOI-6008 and TOI-6086 to exclude the possibility of any background unresolved stellar companion that could be blended with targets at these current positions. Such objects might be introducing the transits signals that we detected in our observations, and skew the results that we obtained from the global analysis.
TOI-5720 has a high proper motion of 350.4 $mas/yr$. We obtained images from POSS-II/DSS (Minkowski & Abell 1963) on 1995, POSS-II/DSS (Lasker et al. 1996) on 1999 and PanSTARRS-1 (Chambers et al. 2016) on 2011 in the blue, infrared and $z_{s}$ bands, respectively, and spanning 28 years with our current observations. TOI-6008 has a relative low proper motion of 158.4 $mas/yr$. We obtained images from POSS-II/DSS (on 1989 in blue band), POSS-II/DSS (on 1994 in the infrared band) and PanSTARRS-1 (on 2011 in the $z_{s}$ band), and spanning 34 years with our current observations. TOI-6086 has also a relative low proper motion of 110.6 $mas/yr$. We also obtained images from POSS-I/DSS (on 1952 in red band), POSS-II/DSS (on 1995 in the infrared band) and PanSTARRS-1 (on 2011 in the $z_{s}$ band), and spanning 71 years with our current observations. TOI-5720, TOI-6008 and TOI-6086 have moved only by $\sim 9.8$, $\sim 5.4$ and $\sim 7.8$ arcsec between 1995 and 2023, 1989 and 2023, and 1952 and 2023, respectively. There is no source in the current day positions of TOI-5720, TOI-6008 and TOI-6086 in any of these archival images (see Figure 9).
Refer to captionRefer to captionRefer to caption
Figure 9: Archive images of TOI-5720 (top panel), TOI-6008 (middle panel) and TOI-6086 (bottom panel). The previous and current positions of the targets are shows in red and cyan colors, respectively. Top row (TOI-5720): 1995 blue image from POSS-II/DSS, 1999 infrared image from POSS-II/DSS2, 2011 $z_{s}$ image from PanSTARRS1, and 2023 $z_{s}$ image from MuSCAT3 (shown from left to right). Middle row (TOI-6008): 1989 blue image from POSS-II/DSS, 1994 infrared image from POSS-II/DSS2, 2011 $z_{s}$ image from PanSTARRS1, and 2023 $z_{s}$ image from MuSCAT3 (shown from left to right). Bottom row (TOI-6086): 1952 red image from POSS-I/DSS, 1995 infrared image from POSS-II/DSS2, 2011 $z_{s}$ image from PanSTARRS1, and 2023 $z_{s}$ image from MuSCAT3 (shown from left to right).
4.5 Statistical validation
We also employed the TRICERATOPS pipeline (Giacalone et al. 2021) to assess the likelihood of false positive scenarios. TRICERATOPS calculates the probability that a light curve originates from a transiting planet or a wide range of other scenarios in a Bayesian framework. In particular, it returns the False Positive Probability (FPP), which is the probability that the observed transit is not caused by a planet (e.g., an eclipsing binary) and the Nearby False Positive Probability (NFPP), which is the probability that the signal comes from a nearby resolved star. Moreover, TRICERATOPS also allows to implement, if available, the high-contrast imaging observations in order to better constrain the calculation. Typically, a planet candidate is considered statistically validated if it returns a FPP$<0.01$ and a NFPP $<0.001$ (Morton 2012, 2015; Giacalone et al. 2021). We ran TRICERATOPS using the light curves of TOI 5720.01, TOI 6008.01 and TOI 6086.01 phase-folded on the orbital periods obtained by the photometric modelling, as we will show in Sect.6. We constraint the overall calculation by implementing the high-contrast imaging observations mentioned in Sect. 4.4. Using ground-based observations (see Sect.4.6), we detected the source of a transit signal on-target, and thus we excluded other nearby sources, which are obviously not contaminating the transit signal, from our FPP calculations, and that is the reason we obtained NFPP $=0$ for the three candidates. We obtained FPP $=(9.6\pm 0.4)\times 10^{-3}$ for the signal associated with TOI 5720.01, FPP $=(0.2\pm 0.05)\times 10^{-3}$ for TOI 6008.01 and FPP $=(5.00\pm 0.05)\times 10^{-3}$ for TOI 6086.01.
4.6 Ground-based follow-up photometric validation
The ground-based photometric follow-up was carried out on the basis of two objectives: i) identify the source of the transit event; ii) chromaticity check for the transit depth in different wavelengths. The closest neighboring stars to TOI-5720 is TIC 903650250 at $54.7\arcsec$ with a $T_{\rm mag}$ of 18.16, and a $\Delta T_{\rm mag}$ of 6.56. The closest neighboring stars to TOI-6008 are respectively TIC 286201110 at $7.3\arcsec$ with a $T_{\rm mag}$ of 17.32, and a $\Delta T_{\rm mag}$ of 5.02, TIC 286201086 at $12.4\arcsec$ with a $T_{\rm mag}$ of 17.43, and a $\Delta T_{\rm mag}$ of 5.13, TIC 286201079 at $14.5\arcsec$ with a $T_{\rm mag}$ of 16.98, and a $\Delta T_{\rm mag}$ of 4.68, and TIC 286201077 at $16.7\arcsec$ with a $T_{\rm mag}$ of 17.59, and a $\Delta T_{\rm mag}$ of 5.29. The closest neighboring stars to TOI-6086 are respectively TIC 18318294 at $25.3\arcsec$ with a $T_{\rm mag}$ of 16.23, and a $\Delta T_{\rm mag}$ of 3.83, and TIC 18318295 at $33.0\arcsec$ with a $T_{\rm mag}$ of 16.81, and a $\Delta T_{\rm mag}$ of 4.41. See the TESS FOV in Figure 1.
We used uncontaminated photometric aperture sizes of only a few arc-seconds using ground-based telescopes, in order to confirm the transit events on the expected target stars for TOI-5720 b, TOI-6008 b and TOI-6086 b. We collected photometric observations for all targets in different bands (zs, Sloan-i’, Sloan-r’, Sloan-g’), spanning from 477 to 1100 nm. We measured a matching transit depth within $1\sigma$ in all bands. The transit depths obtained in different bands for TOI-5720 b, TOI-6008 b, and TOI-6086 b are presented in Figure 10 and Table 4.
Figure 10: Measured transit depths in different filters (colored dots with error bars) obtained in the global analysis for TOI-5720 b (top panel), TOI-6008 b (middle panel) and TOI-6086 b (bottom panel). The dashed line corresponds to the depth obtained from the achromatic fit in the global analysis with $1\sigma$ error bar (shaded region). All measurements agree with the common transit depth at $1\sigma$.
5 Planet searches using the TESS photometry
Using the TESS data described in Section 2.1 and our SHERLOCK package (Pozuelos et al. 2020; Demory et al. 2020) we aimed to (i) recover the candidates TOI-5720.01, TOI-6008.01, and TOI-6086.01 and (ii) search for extra transiting planets in these systems that may remain unnoticed due to detection thresholds of the SPOC. The SHERLOCK pipeline is specially set up for searching low-S/N transit-like features that might be attributable to planets. This pipeline is used to explore space-based data produced by the TESS and Kepler/K2 missions; for example, it is used in the SPECULOOS project (Sebastian et al. 2021), where is searching for transiting planets orbiting ultra-cool dwarfs and the FATE project (Van Grootel et al. 2021), where is searching for planetary remnants orbiting hot subdwarfs (Thuillier et al. 2022). In particular, it uses a multi-detrend approach applying a bi-weight filter several times with different window sizes and the TLS algorithm (Hippke & Heller 2019) to optimize the transit search by converging to the signals with the highest S/N and SDE (signal-detection-efficiency) (see, e.g., Delrez et al. 2022; Pozuelos et al. 2023). In all the cases, we first found the signal corresponding to the planetary candidate alerted by SPOC; that is, we successfully recovered all the TOIs studied in this work. In subsequent runs, for TOI-5720 and TOI-6086, we found no other signal that might be attributable to planets. On the other hand, for TOI-6008, we found a weak signal corresponding to an Earth-sized planet of $\sim$1.12 R with an orbital period of 16.35 days. The semiautomatic vetting conducted by SHERLOCK did not find any hint of systematics that might be producing the signal; however, the statistical validation (automatically carried out by SHERLOCK but relaying in TRICERATOPS) yielded an NFPP and FPP of 0.175 and 0.356, placing it as a likely false positive (Giacalone et al. 2021). Still, we decided to conduct ground-based observations to fully confirm this signal’s false positive character. We conducted three observations with MuSCAT2 on August 03 2023 and one with MuSCAT3 on August 19 2023; both of the cases, we do not find any evidence of a transit feature in the data, allowing us to conclude that this signal is indeed a false positive. Moreover, if the planet was present, we would see a clear transit in the MuSCAT3 light curve (See Figure 11) with a signal-to-noise ratio of $S/N\sim 6$. Finally, we computed also the Lomb–Scargle periodogram (Lomb 1976; Scargle 1982), there were no indications of flaring activity or stellar modulation seen in the TESS data of these three systems. This implies that the rotational periods of stars are probably longer than the TESS observation window for a single sector.
Figure 11: LCO-HAL-2m0/MuSCAT photometry of TOI-6008.22 collected on UT August 20, 2023in the Sloan-$i^{\prime}$ filter with an exposure time of 30s. The blue data points show the raw light curve. The red data points show the light curve after the transit injection of an Earth-sized planet of $\sim 1.12$ R with the best-fitting model is superimposed in black. The green region shows the predicted ingress and egress of the transit. The light curves are shifted along the y-axis for clarity.
6 Global modelling: Photometric and RVs
We performed a global fit of all available transit light curves from TESS, MuSCAT2 and MuSCAT3 telescopes, and radial velocity measurements from Subaru-8.2m/IRD spectrograph described in Sections 2, using the Metropolis-Hastings (Metropolis et al. 1953; Hastings 1970) method implemented in TRAFIT, a revised version of Markov-chain Monte Carlo (MCMC) code described in Gillon et al. (2010, 2012, 2014). The RVs collected by Subaru/IRD spectrograph were modeled with a two-body Keplerian model (Murray & Correia 2010), while the transit light curves were modeled using the Mandel & Agol (2002) quadratic limb-darkening model, multiplied by a transit baseline, aiming to correct for external systematic effects (time, airmass, x- and y- position of the star on the detector, background and FWHM of the PSF).
The baseline model for each transit was selected by minimizing the BIC (Bayesian information criterion, Schwarz (1978)), and photometric error bars were re-scaled using the correction factor $CF=\beta_{w}\times\beta_{r}$, where, $\beta_{r}$ is the red noise and $\beta_{w}$ is the white noise (Gillon et al. 2012). The error bars of TOI-6008 RVs measurements were quadratically re-scaled using the ”jitter” noise $=(<Error_{RVs}>^{2}-\sigma^{2}_{RVs})^{1/2}$, where, $<Error_{RVs}>$ is the mean error of the RVs measurements, and $\sigma_{RVs}$ is the standard deviation of the best-fit residuals.
For each system, the jump parameters that were sampled by MCMC: the transit epoch, the planet orbital period, impact parameter, the transit depth, the total transit duration, and the combination $q_{1}=(u_{1}+u_{2})^{2}$ and $q_{2}=0.5u_{1}(u_{1}+u_{2})^{-1}$ (Kipping 2013) of its quadratic limb-darkening coefficients $u_{1}$ $u_{2}$, which are calculated from Claret et al. (2012); Claret (2018).
For each target, we applied a Gaussian prior distribution on the stellar parameters derived from the SED and spectroscopy analysis ($R_{\star}$: stellar radius, $M_{\star}$: stellar mass, $T_{\rm eff}$: stellar effective temperature, $[Fe/H]$: stellar metallicity and $\log g_{\star}$: stellar surface gravity). Two MCMC analyses were performed for each system, the first assuming a circular orbit (i.e $ecc=0$) and the second assuming an eccentric orbit. Based on the Bayes factor $BC=\exp{(-\Delta BIC/2)}$, our results are well compatible with a circular orbit solution.
For each transit light curve, we performed a preliminary analysis composed of 1$\times$×Markov-chain of 100,000 steps to compute the $CF$. The second step, we performed a global MCMC analysis composed of 3$\times$ Markov-chains of 100,000 steps to infer the stellar and planetary physical parameters for each system. We used the Gelman & Rubin (1992) statistical test in order to check for the convergence of each Markov chain.
The results of TOI-5720, TOI-6008 and TOI-6086 are presented in Tables 3, 4 and 5.
7 Results and discussion
We present the validation and discovery of TOI-5720 b, TOI-6008 b and TOI-6086 b by the TESS mission, which were confirmed using follow-up photometric observations obtained by MuSCAT2, MuSCAT3 and LCOGT-1m0 telescopes (see Figure 3), and radial velocity measurements obtained by Subaru-8.2m/IRD spectrograph (see RVs curve in Figure 6). We characterized the host-stars by combining spectroscopic observations obtained with Shane/Kast (see Section 2.2.4) and IRTF/SpeX (see Section 2.2.5), spectral energy distribution (SED), and stellar evolutionary models (see Section 3.3). We performed a global analysis of TESS data, and ground-based photometric and radial velocity observations in order to constrain the physical parameters for each system (see Section 6). Table 3 shows the stellar properties of host stars TOI-5720, TOI-6008 and TOI-6086. Derived stellar and planetary physical parameters for each system are shown in Tables 3, 5 and 4. Figures 19, 20 and 21 show the parameters posterior distributions for TOI-5720, TOI-6008 and TOI-6086, respectively.
7.1 TOI-5720 b, TOI-6008 b and TOI-6086 b
The host star TOI-5720 is a nearby (23 pc) $K_{\rm mag}=9.54$ M3.5-type star, with an effective temperature of $T_{\rm eff}=3325\pm 75$ K, a surface gravity of $\log g_{\star}=5.01\pm 0.02$ dex, a radius of $R_{\star}=0.383\pm 0.018\leavevmode\nobreak\ R_{\odot}$ and a mass of $M_{\star}=0.383\pm 0.019\leavevmode\nobreak\ M_{\odot}$ (derived from the SED analysis), and a metallicity of $[Fe/H]=-0.19\pm 0.12$ (deduced from the near-infrared spectrum). TOI-5720 b is a short-period Earth-sized planet orbiting around its host star every $P=1.4344555\pm 0.0000036$ days, which has a radius of $R_{p}=1.09\pm 0.07\leavevmode\nobreak\ R_{\oplus}$, an equilibrium temperature of $T_{\rm eq}=708\pm 19$ K (assuming an albedo of 0), an incident flux of $S_{p}=41.7\pm 4.5$ times that of Earth, and an upper limit on the mass of $M_{p}<4\leavevmode\nobreak\ M_{\oplus}$ based on 2 TRES RV measurements with $3\sigma$ error bar (see Section 2.3.2), this confirmed the non-stellar transiting companion. Moreover, the predicted mass using the Chen & Kipping (2017) mass-radius relationship implemented in the forecaster555forecaster: https://github.com/chenjj2/forecaster package, it is found to be $M_{p}=1.32^{+0.92}_{-0.47}$ M.
TOI-6008 is an M5-type ($K_{\rm mag}=9.29$) star at a distance of 36 pc, with an effective temperature of $T_{\rm eff}=3075\pm 75$ K, a surface gravity of $\log g_{\star}=5.01\pm 0.04$ dex, a radius of $R_{\star}=0.242\pm 0.013\leavevmode\nobreak\ R_{\odot}$ and a mass of $M_{\star}=0.230\pm 0.011\leavevmode\nobreak\ M_{\odot}$ (derived from the SED analysis), and a metallicity of $[Fe/H]=-0.21\pm 0.12$ (deduced from the near-infrared spectrum). TOI-6008 b is a $1.03\pm 0.05\leavevmode\nobreak\ R_{\oplus}$ short-period planet with an orbital period of $P=0.8574347\pm 0.0000424$ day, an equilibrium temperature of $T_{\rm eq}=707\pm 19$ K (assuming an albedo of 0) and an incident flux of $S_{p}=41.5\pm 4.5$ times that of Earth. We collected 15 RV measurements of TOI-6008 using the Subaru/IRD spectrograph, the resulting mass is found to be $M_{p}=1.1^{+3.3}_{-0.9}\leavevmode\nobreak\ M_{\oplus}$ with $3\sigma$ error bar. This value ruled out stellar transiting companion, and confirmed a planetary companion nature. Additionally, the predicted planetary mass using the mass-radius relationship implemented in the forecaster package, the value is found to be $M_{p}=1.10^{+0.73}_{-0.36}$ M.
TOI-6086 is a $K_{\rm mag}=9.99$ M3-type star at a distance of 32 pc from the Sun, with an effective temperature of $T_{\rm eff}=3200\pm 75$ K, a surface gravity of $\log g_{\star}=5.01\pm 0.02$ dex, a radius of $R_{\star}=0.259\pm 0.013\leavevmode\nobreak\ R_{\odot}$ and a mass of $M_{\star}=0.254\pm 0.011\leavevmode\nobreak\ M_{\odot}$ (derived from the SED analysis), and a metallicity of $[Fe/H]=-0.2\pm 0.3$ (deduced from the near-infrared spectrum). TOI-6086 b is a short-period ($P=1.3888725\pm 0.0000827$ days) Earth-sized planet ($R_{p}=1.18\pm 0.07\leavevmode\nobreak\ R_{\oplus}$) orbiting around its host star, with an equilibrium temperature of $T_{\rm eq}=634\pm 16$ K (assuming an albedo of 0), an incident flux of $S_{p}=26.8\pm 2.7$ times that of Earth. Using the mass-radius relationship implemented in the forecaster package, we find that TOI-6086 b has a predicted mass of $M_{p}=1.68^{+1.20}_{-0.55}$ M.
Figures 12 and 13 show the distribution of our planets versus orbital period and equilibrium temperature. These systems are suitable targets for high precision radial velocity facilities in order to provide the planetary mass measurements, bulk densities as well as to search for additional outer planets in the system. Assuming a rocky composition for these planets, the predicted RV semi-amplitudes using the mass-radius relationship by Chen & Kipping (2017), the values are found to be $K_{\rm TOI-5720}=1.4^{+1.0}_{-0.5}$ m/s for TOI-5720, $K_{\rm TOI-6008}=2.0^{+1.3}_{-0.6}$ m/s for TOI-6008 and $K_{\rm TOI-6086}=2.4^{+1.6}_{-0.8}$ m/s for TOI-6086. By combining the predicted RV semi-amplitudes with the relative faintness of the host stars ($V_{\rm mag}$ of 13.9 to 16.0), would require a high precision spectroscopic observations using stabilized spectrographs such as MAROON-X (Seifahrt et al. 2018) mounted on the 8.1m-Gemini-North telescope. Assuming an exposure times of 1800s and good observing conditions, the MAROON-X exposure time calculator (ETC) predict RV precision of 0.6 m/s, 1.1 m/s and 1.0 m/s for TOI-5720, TOI-6008 and TOI-6086, respectively. Using a few RV measurements for each target will allow us to measure the planetary masses, bulk densities and the eccentricities of the orbits with 10-$\sigma$ precision.
The combination of the equilibrium temperature, the large stellar incident flux, the infrared brightness and small sizes of the host stars, making all three planets suitable for transmission and emission spectroscopy studies with JWST. We calculated the transmission spectroscopy metrics (TSM) using equation #1 from Kempton et al. (2018). Our predicted TSM are found to be $9.7^{+7.1}_{-4.1}$ for TOI-5720 b, $18.8^{+12.6}_{-6.8}$ for TOI-6008 b and $12.0^{+8.5}_{-4.7}$ for TOI-6086 b. We also calculated their emission spectroscopy metrics (ESM) using equation #4 from Kempton et al. (2018), and we find that ESM to be $3.4\pm 0.5$ for TOI-5720 b, $6.4\pm 0.7$ for TOI-6008 b and $4.2\pm 0.7$ for TOI-6086 b. Figure 14 shows the planets positions in the TSM and ESM vs planetary radius diagrams of Earth-sized ($R_{p}<1.5$ R) transiting exoplanets around M-dwarfs ($T_{\rm eff}<3800$) observed or scheduled on JWST. Based of TSM and ESM parameters, TOI-6008 b stands out as one of the favourable terrestrial planet for atmospheric characterization (emission, transmission and phase curve photometry) compared to other systems that have been observed with JWST (e.g. TRAPPIST-1 b, TRAPPIST-1 c, TRAPPIST-1 e and TRAPPIST-1 g; Gillon et al. (2017), L 98-59 b; Kostov et al. (2019), GJ 357 b; Luque et al. (2019), GJ 3473 b; Kemmer et al. (2020), LHS 1140 c; Dittmann et al. (2017), LTT 1445A b; Winters et al. (2019), LTT 1445A c; Winters et al. (2022), TOI-1468 b; Chaturvedi et al. (2022), LHS 1478 b; Soto et al. (2021), TOI-732 b; Nowak et al. (2020), LP 791-18 b; Crossfield et al. (2019)). Thanks to the relative high equilibrium temperatures of the planets, small sizes and the brightness of the stars, making also TOI-5720 b and TOI-6086 b good targets for the emission spectroscopy and phase curve photometry compared with TRAPPIST-1 b and TRAPPIST-1 c.
Figure 12: Planetary radius as a function of the orbital period diagram of known transiting exoplanets from NASA Archive of Exoplanets. Grey circles show the FGK planetary systems, and colored points show M-dwarfs systems. M-dwarfs planets are colored by the host star effective temperature. The size of the points scale according to the stellar radius. All short-period ($<$1.5 days) Earth-sized planets orbiting around M-dwarfs are highlighted with error bars, and planets with and without mass measurements are highlighted with + and -, respectively. TOI-5720 b, TOI-6008 b and TOI-6086 b are circled in blue, purple and green, respectively.
Figure 13: Planetary radius as a function of planetary equilibrium temperature diagram of known transiting exoplanets from NASA Archive of Exoplanets. Grey circles show the FGK planetary systems, and colored points show M-dwarfs systems. M-dwarfs planets are colored by the host star effective temperature. The size of the points scale according to the stellar radius. TOI-5720 b, TOI-6008 b and TOI-6086 b are circled in blue, purple and green, respectively.
7.2 Prospects for atmospheric characterization
Here we provide a thorough examination of the simulated transmission spectra of planets utilizing PLanetary Atmospheric Transmission for Observer Noobs (PLATON; Zhang et al. 2019, 2020). Specifically, we employed PandExo (Batalha et al. 2017) to generate synthetic data, mimicking observations that would be made with the James Webb Space Telescope (JWST). Our simulation methodology encompasses various parameters such as planetary and stellar radii, effective temperatures, stellar metallicity, stellar gravity, transit duration, orbital period, and inclination. Through meticulous adjustment of these inputs, we created mock datasets to replicate transmission spectra. We ensured a minimum of $3$ transit observations, spanning between $3$-$4$ hours of observation time, with at least one-half of the time allocated for out-of-transit observations for each target to acquire the necessary simulated data. The resultant synthetic transmission spectra offer insights into the atmospheric properties of these planets. Through detailed simulation and analysis of these spectra, our goal is to contribute valuable information regarding the composition and characteristics of exoplanetary atmospheres. Figure 15 presents the synthetic spectra for TOI-5720 b, TOI-6008 b and TOI-6086 b.
The James Webb Space Telescope (JWST) is equipped with four scientific instruments (Gardner et al. 2006), delivering high-resolution and highly sensitive observations of celestial targets. The JWST NIRSpec (Rauscher et al. 2003) employs fixed slits (FS) spectroscopy, offering high sensitivity single-object spectroscopy across a wide wavelength range from 0.6 $\mu$m to 5.3 $\mu$m. The JWST NIRCam (Horner et al. 2002), in its wide-field slitless spectroscopy mode, utilizes grisms and filters covering wavelengths between 2.4 $\mu$m to 5.0 $\mu$m, achieving a resolving power of $1,600$ at 4 $\mu$m. Our study utilized the F322W2 and F [^4] W filters, spanning wavelengths from 2.7 $\mu$m to 4 $\mu$m and 4 $\mu$m to 5 $\mu$m, respectively. The JWST Near Infrared Imager and Slitless Spectrograph (NIRISS; (Doyon & Hutchings 2012)) provide slit-less spectroscopy across wavelengths ranging from 0.6 $\mu$m to 5.0 $\mu$m. The Mid-Infrared Instrument (MIRI; (Boccaletti et al. 2003)) extends our observational capabilities into the mid-infrared spectrum, covering spectral wavelengths from 4.9 $\mu$m to 28.8 $\mu$m.
In our investigation of three validated planets, we leverage various grism and filter combinations available for each instrument. These combinations act as instrumental tools facilitating a comprehensive simulation of the transmission spectra for each planet.
7.2.1 PLATON
PLATON (Zhang et al. 2019, 2020), a Python package, computes transmission spectra for exoplanets and derives atmospheric characteristics based on observed spectra. The software accommodates key atmospheric parameters, including temperature, metallicity, C/O ratio, cloud-top pressure, and scattering slope. Employing an equilibrium chemistry model computed with GGchem (Woitke et al. 2018), PLATON allows users to input metallicity and C/O ratio. From these inputs, PLATON calculates the abundances of $34$ chemical species in each atmospheric layer. GGchem computes these abundances across a grid of metallicity, C/O ratio, temperature, and pressure, assuming equilibrium chemistry with or without condensation. The parameter ranges are as follows: metallicity from $0.1$ to $1000$ times solar, C/O ratio from $0.2$ to $2$, temperature from $300$ to $3000$ K, and pressure from $10^{-4}$ to $10^{8}$ Pa. To compute a transmission spectrum, we adopt the assumption of an isothermal atmosphere with equilibrium chemistry. PLATON divides the atmosphere into 500 equally spaced layers in $\log{P}$ from $10^{-4}$ to $10^{8}$. The physical depth of each layer is determined by applying the hydrostatic equation. The opacity of each layer is subsequently calculated, taking into account factors such as gas absorption, collisional absorption, and Rayleigh scattering. The radiative transfer calculation is executed by considering a ray passing through the atmosphere with an impact parameter. The optical depth experienced by the ray is then computed, and the transit depth is derived by summing the cross-sectional areas of each layer, with weights assigned based on the amount of light permitted through. Subsequently, corrections for unocculted starspots are implemented, which are deemed crucial for active stars due to the distinct spectra of both hot, unspotted regions and colder starspots. The correction involves adjusting the transit depth based on factors such as the spot fraction, spot temperature, temperature of unspotted regions, and the spectrum of the stellar surface. Each layer’s composition is determined by considering equilibrium chemistry, which facilitates the calculation of the mean molecular weight. Fitted parameters include the mixing ratios of these molecules, the atmospheric mean molecular weight, the surface pressure, and the planet’s radius. The specified radius for the planet is taken at a reference pressure of $1$ bar, chosen for its proximity to the standard atmospheric pressure on Earth. Our model assumes a clear sky without clouds or haze.
7.2.2 PandExo
PandExo (Batalha et al. 2017) serves as both an online interface and an open-source Python package, employed in modeling instrumental noise for exoplanet transit time series spectroscopy modes with JWST (NIRISS, NIRCam, NIRSpec, and MIRI LRS) and HST (WFC3). Drawing on the foundations of the Space Telescope Science Institute’s (STScI) Exposure Time Calculator, Pandeia (Pontoppidan et al. 2016), PandExo exhibits excellent agreement, with deviations well within $10\%$ not only for HST WFC3 but also across the entire spectrum of JWST instruments. The input parameters encompass critical factors, including a stellar Spectral Energy Distribution (SED) model sourced from the Phoenix Stellar Atlas (Husser et al. 2013b), the apparent magnitude of the star, details about the planet’s spectrum (primary or secondary), transit duration, the ratio of time dedicated to in-transit versus out-of-transit observations, the number of observed transits, the saturation threshold expressed as a percentage of the full well capacity, and the user-defined noise floor. For the computation of the planet’s atmospheric transmission data, we utilized the built-in molecular lines grid in PandExo. Our approach involved selecting the planetary temperature closest to the computed temperature of each of the eight validated planets. We considered equilibrium chemistry with TiO and medium Rayleigh Scattering. It is essential to note that our modeling assumes conditions devoid of clouds and hazes, representing clear skies for these analyses.
Refer to captionRefer to caption
Figure 14: Feasibility of TOI-5720 b, TOI-6008 b and TOI-6086 b for transmission and emission spectroscopy studies. Transmission spectroscopy metric (TSM, top panel) and emission spectroscopy metric (ESM, bottom panel) as a function of planetary radius. TSM and ESM are computed using equations 1 and 4 from Kempton et al. (2018). Our planets are compared to the Earth-sized ($R_{p}<1.5$ R) known transiting exoplanets observed or scheduled on JWST around M-dwarfs ($T_{\rm eff}<3800$). The size of the points scale according to the stellar radius. The points are coloured according to the stellar effective temperature. The blue circles highlight our planets TOI-5720 b, TOI-6008 b and TOI-6086 b.
7.3 Results
In this section we discuss about the synthetic spectra we obtained using PandExo and best fit model we find using PLATON. In Table 6 we provided the details of exposure time and SNR values.
7.3.1 TOI-5720 b
Our model spectra using PLATON align well with the simulated data from PandExo. $\rm H_{2}O$ lines can be found at 1, 1.3, 1.8, and 2.8 $\mu$m using NIRSpec, NIRISS, and NIRCam. Observation using MIRI may also add a contribution to detecting subtle $\rm H_{2}O$ signatures at higher wavelengths. We also found $\rm CO_{2}$ signature at 4.3 $\mu$m, depending on the metallicity and C/O chosen for our study. Our fitting efforts yield a best-fit model with an estimated metallicity of the atmosphere of TOI-5720b to be 1.26 times Solar metallicity and C/O to be 1.01. We also find $\rm CH_{4}$ lines at 1.2, 1.4, 1.7, 2.3, and 3.3 $\mu$m with NIRSpec, NIRISS, and NIRCam (see top panel of Figure 15). There are also a possible presence of $\rm NH_{3}$ in small amount.
7.3.2 TOI-6008 b
Our model spectra using PLATON align well with the simulated data from PandExo. $\rm H_{2}O$ lines can be found at 0.8, 1.2, 1.4, 1.8, and 2.7 $\mu$m using NIRSpec, NIRISS, and NIRCam. Observation using MIRI may also add a contribution to detecting subtle $\rm H_{2}O$ signatures at higher wavelengths. We also found $\rm CO_{2}$ signature at 4.3 $\mu$m and $\rm CO$ at 4.8 $\mu$m, depending on the metallicity and C/O chosen for our study. Our fitting efforts yield a best-fit model with an estimated metallicity of the atmosphere of TOI-6008b to be 1.26 times Solar metallicity and C/O to be 0.5 (see middle panel of Figure 15).
7.3.3 TOI-6086 b
Our model spectra using PLATON align well with the simulated data from PandExo. $\rm H_{2}O$ lines can be found at about 0.8, 1.0, 1.3, 1.8, and 3 $\mu$m using NIRSpec, NIRISS, and NIRCam. Observation using MIRI may also add a contribution to detecting subtle $\rm H_{2}O$ signatures at higher wavelengths. We also found $\rm CO_{2}$ signature at 4.3 $\mu$m and $\rm CO$ at 4.8 $\mu$m, depending on the metallicity and C/O chosen for our study. Our fitting efforts yield a best-fit model with an estimated metallicity of the atmosphere of TOI-6086b to be 1.26 times Solar metallicity and C/O to be 0.3. There is also a possible small amount of $\rm NH_{3}$ present. We also find $\rm CH_{4}$ lines at 0.8, 1.2, 1.4, 1.7, 2.3, and 3.3 $\mu$m with NIRSpec, NIRISS, and NIRCam (see bottom panel of Figure 15).
Figure 15: Synthetic transmission spectrum in JWST for TOI-5720 b (top panel), TOI-6008 b (middle panel) and TOI-6086 b (bottom panel) generated with the help of PLATON and PandExo tools. Red, yellow, green and grey points are simulated data using JWST/MIRI, JWST/NIRSpec, JWST/NIRCam and JWST/NIRISS instruments, and the best-fitting model is superimposed in blue.
Instrument Mode a Exposure SNR
Time
TOI 5720 b
MIRI Slit-less Spectroscopy 7003 s 3095.56
NIRSpec G140H/F070LP 7179 s 3906.26
G140H/F100LP 7168 s 3713.27
G235H/F170LP 7170 s 4072.01
G395H/F290LP 7177 s 3693.03
NIRCam F322W2 7221 s 3514.38
F [^4] W 7177 s 2894.01
NIRISS SUBSTRIP256 7211 s 5645.12
TOI 6008 b
MIRI Slit-less Spectroscopy 8370 s 2513.81
NIRSpec G140H/F070LP b 8526 s 3744.90
G140H/F100LP 8510 s 4142.91
G235H/F170LP 8534 s 4258.71
G395H/F290LP 8518 s 3555.67
NIRCam F322W2 8635 s 3536.38
F [^4] W 8500 s 2755.55
NIRISS SUBSTRIP96 c 8664 s 6833.68
TOI 6086 b
MIRI Slit-less Spectroscopy 8222 s 1934.40
NIRSpec G140H/F070LP 8303 s 3115.41
G140H/F100LP 8315 s 3345.80
G235H/F170LP 8350 s 3214.50
G395H/F290LP 8376 s 3037.43
NIRCam F322W2 8431 s 2903.43
F [^4] W 8516 s 2342.96
NIRISS SUBSTRIP256 8386 s 5107.16
a Details about different modes, resolving power and wavelength can be found on: https://jwst-docs.stsci.edu/
b Disperser-filter combinations. We chose combinations that can provide a high resolution power of   2700. G = Grism, F = Filter.
c Size of sub array for single object slitless spectroscopy.
Table 6: Exposure times and SNR obtained for each synthetic specrum of JWST.
8 Software
This work made use of many publicly available packages & tools for which the authors are immensely grateful. In addition to the software and pipelines cited throughout the paper, we also used Matplotlib (Hunter 2007), Numpy (van der Walt et al. 2011), Astropy (Astropy Collaboration et al. 2013), SciPy (Jones et al. 2001), Wotan (Hippke et al. 2019), and Photutils (Bradley et al. 2020).
9 Acknowledgements
The postdoctoral fellowship of KB is funded by F.R.S.-FNRS grant T.0109.20 and by the Francqui Foundation. This paper made use of data collected by the TESS mission and are publicly available from the Mikulski Archive for Space Telescopes (MAST) operated by the Space Telescope Science Institute (STScI). Funding for the TESS mission is provided by NASA’s Science Mission Directorate. We acknowledge the use of public TESS data from pipelines at the TESS Science Office and at the TESS Science Processing Operations Center. Resources supporting this work were provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center for the production of the SPOC data products. This work is partly supported by JSPS KAKENHI Grant Numbers JP18H05439, JP 18H05442, JP21K20376, and JST CREST Grant Number JPMJCR1761. This article is based on observations made with the MuSCAT2 instrument, developed by ABC, at Telescopio Carlos Sánchez operated on the island of Tenerife by the IAC in the Spanish Observatorio del Teide. This paper is based on observations made with the MuSCAT3 instrument, developed by the Astrobiology Center and under financial supports by JSPS KAKENHI (JP18H05439) and JST PRESTO (JPMJPR1775), at Faulkes Telescope North on Maui, HI, operated by the Las Cumbres Observatory. (Some of the) Observations in the paper made use of the NN-EXPLORE Exoplanet and Stellar Speckle Imager (NESSI). NESSI was funded by the NASA Exoplanet Exploration Program and the NASA Ames Research Center. NESSI was built at the Ames Research Center by Steve B. Howell, Nic Scott, Elliott P. Horch, and Emmett Quigley. We acknowledge financial support from the Agencia Estatal de Investigación of the Ministerio de Ciencia e Innovación MCIN/AEI/10.13039/501100011033 and the ERDF “A way of making Europe” through project PID2021-125627OB-C32, and from the Centre of Excellence “Severo Ochoa” award to the Instituto de Astrofisica de Canarias. This work makes use of observations from the LCOGT network. Part of the LCOGT telescope time was granted by NOIRLab through the Mid-Scale Innovations Program (MSIP). MSIP is funded by NSF. This research has made use of the Exoplanet Follow-up Observation Program (ExoFOP; DOI: 10.26134/ExoFOP5) website, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program. MG is F.R.S.-FNRS Research Director. KAC and CNW acknowledge support from the TESS mission via subaward s3449 from MIT. B.V.R. thanks the Heising-Simons Foundation for Support. This material is based upon work supported by the National Aeronautics and Space Administration under Agreement No. 80NSSC21K0593 for the program “Alien Earths”. The results reported herein benefited from collaborations and/or information exchange within NASA’s Nexus for Exoplanet System Science (NExSS) research coordination network sponsored by NASA’s Science Mission Directorate. F.J.P acknowledges financial support from the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/10.13039/501100011033 and through the project PID2022-137241NB-C43.
Appendix A Discovery TESS light curves for TOI-5720, TOI-6008 and TOI-6086
Figure 16: TESS  phase-folded photometric data of TOI-5720 b (blue), TOI-6008 b (purple) and TOI-6086 b (green), with best-fitting transit model is superimposed.
Appendix B High-Resolution imaging: WIYN 3.5 m/NESSI
We observed TOI-5720 on UT 2023 February 5 using the NN-EXPLORE Exoplanet Stellar Speckle Imager (NESSI; Scott et al. (2018)), a speckle imager employed at the WIYN 3.5 m telescope on Kitt Peak. NESSI was used to obtain simultaneous speckle imaging in two filters with central wavelengths $\lambda_{c}=562$ and 832 nm. The observation consisted of a set of 9 1000-frame 40 ms exposures. NESSI’s field-of-view was set by a $256\times 256$ pixel sub-array readout, which resulted in a $4.6\times 4.6$ arcsecond field. However, our speckle measurements were confined to an outer radius of 1.2 arcseconds from the target star. Speckle imaging of a point source standard star was also taken in conjunction to the observation of the TOI. The standard observation consisted of a single 1000-frame image set and was used to calibrate the underlying PSF.
These speckle data were reduced using a pipeline described by Howell et al. (2011). Among the pipeline products is a reconstructed image of the field around TOI-5720 in each filter. The reconstructed image was used to measure a contrast curve setting the detection limits on any point sources in close proximity to TOI-5720. The contrast level was obtained by measuring fluctuations in the noise-like background level as a function of separation from the target star. No companion sources were detected for TOI-5720 in the NESSI data (see Figure 17).
Figure 17: High resolution imaging of TOI-5720 obtained with the WIYN-3.5m/NESSI on UT 2023 Feb 5 in two filters with central wavelengths $\lambda_{c}=562$ and 832 nm. No close companion sources were detected near the target.
Appendix C High-Resolution imaging: Robo AO
We used additional sub-arcsecond imaging taken with Robo-AO, an autonomous laser-guided adaptive optics (AO) system (Baranec et al. 2014). The Robo-AO M-dwarf multiplicity catalog contains AO imaging of TOI-5720 and TOI-6008, taken on 2016 February 18 and 2016 June 18, respectively, with the 2.1m Kitt Peak telescope (Lamman et al. 2020). The images were taken with an Andor iXon DU-888 camera in the $i^{\prime}$ filter with an exposure time of 90s. Median seeing at the telescope was 1.44” which resulted in a FWHM of around 0.12”. Observations were then processed with automatic pipelines, including one optimized for either high or low signal-to-noise images and a ”high-contrast” imaging pipeline which creates a locally-optimized PSF-subtracted image around the target using reference starts observed the same night (Jensen-Clem et al. 2018). Contrast curves of the observations are shown in Figure 18; there is no indication of any additional close stellar companions were detected above these contrast curves.
Figure 18: Robo-AO contrast curves and images of TOI-5720 (top) and TOI-6008 (bottom). These observations were taken via an automated laser-guided adaptive optics system on the 2.1m Kitt Peak telescope. No stellar companions were detected in these observations (Lamman et al. 2020).
Appendix D Posterior probability distribution for the TOI-5720, TOI-6008 and TOI-6086 systems
Figure 19: The PDF (Posterior probability distribution) for the TOI-5720 system parameters derived from our global MCMC analysis. The median value for each parameter is represented by the vertical dashed lines.
Figure 20: The PDF for the TOI-6008 system.
Figure 21: The PDF for the TOI-6086 system.